From GlueXWiki
Jump to: navigation, search

December 6, 2012 FDC meeting


  1. Production Construction Tracking
    • Production status (Dave)
    • Package testing (Lubomir)
  2. Corrosion tests (Lubomir)
  3. Engineering (Bill)
  4. Electronics (Chris, Nick)
  5. Tests at EEL126 FDC E-log
  6. Other


Participants: Eugene, Bill, Dave, Chris, Nick, Simon, Beni, and Lubomir.


- Dave: priority is to finish the two type-3 cathodes so that we can build three cells from package #2. The first type-3 will be ready on Monday, then we can start the installation. Working on all the other cathodes for this package. Since Tina and Mike will start working also on BCAL we'll have to find a way to finish at least the package#2 cathodes, that has to be done by the end of January.

- Package #4: the broken wire (top cell) was replaced and the package put together last Friday. This week the top cell was tested: all channels working fine, very low currents of 15/-4nA, oxygen after six days flushing ~160ppm. The plan is to take the package from the installation table and start putting the other subsystems. The problem is that we still don't have the clips.

- All the subsystems on package #1 were installed.

- Package #3: as explained in entry #98, the package was connected to the Nitrogen dewar with a rotameter at 30sccpm, and there was an over-pressure over the last weekend. We measured the pressure at the output of the rotameter: 360Pa, so the pressure in the chamber must have been lower than that.


- Chris: delay with the clips due to lack of material. They made only 135 (out of 500) and Chris asked them to ship these; plus about 50 here could be enough for one package. Nick populated one more set of PCB and will do another (the last) one soon. Chris and Nick finished the HV cabling on package #1.

- According to Beni, Cody made progress in coding the firmware for fADC125 and will do some tests in 126.

Cathode corrosion test

- Lubomir: test #3 (Ar/CO2 gas, at ~70degC). One more sample failed over the weekend, 25 days after the start of the test: one of the two Viton samples, second Viton is OK. Again we suspect it is the solder that failed.


- Bill ordered fiducial stuff the parts for the package spacers. All the parts for the cooling, except the chiller, are here. Will talk to the survey people about the details.

- Bill complained about many corrupted files on the M-drive in the halld_engineering directory. Most of them have also files with the correct versions, but some, including several FDC procedures, don't.

Tests at EEL126

- The two AC units in 126 are now installed, both the inside and outside parts, but not connected electrically; will happen next week.

- Lubomir showed latest results for the wire resolution using the three-cathode configuration for the third package; details in the FDC Elog #99. After extracting experimentally the time-to-distance function using cathode information only, then the track can be reconstructed with a minimum number of cells. This first step is needed to find the offsets of all the wire planes. Then one can do the track fitting using all the planes except one (cell #2) on which we study the resolution. The residuals for cell #2 showed better resolution (second plot in #99), ~190 microns on average, than obtained before for cell #3. In case of cell #3 one had to take into account the strip resolution, and the value for that was too optimistic, this can explain the cell #2/3 resolution differences. Now the resolution is closer to that obtained with the prototype but without the time-to-distance conversion, except at big distances from the wire. There the field is not homogeneous and the statistics is very poor, both these can explain the problem.

- The offsets of the wires were found w.r.t. frame defined by cells #1 and #4 and they are: -282, -116, +409, +147 microns for cells #2,3,5,6. So, somehow cell #5 is off by 200-300 microns. By design this should not happen if the strong-back doesn't move w.r.t pin-rails. The strong-back is attached to the granite table with micrometers that are locked, but Dave said sometimes these needed to be adjusted. We discussed if the Beni's results about deviations from 60deg cell rotation may have the same origin and related questions about surveying and extracting the offsets from the data.


- Beni is working on the HV GUI for the FDC.

- Next FDC meeting will be January 3, due to the physics Christmas party next Thursday.